Your company is evaluating an investment opportunity that will require an initial cash outlay of $10,000. It is anticipated that the investment will generate $2,000 in added annual revenues for 10 years. The cost of capital for the project is 10%.

Examples of cash flow considerations: (based on the project described above)

For each of the project details provided below, indicate whether

1) It should be included in calculating the company’s NPV for the project, and

2 ) How the data should be included (as part of the initial investment, the project’s cash flows, or interest rate)

a. As a maintenance project, the project’s risk will be 3% lower than the typical investment for the company.

b. The company will need to pay maintenance costs of $500 a year.

c. $5,000 of the initial investment is for a piece of equipment that the company owns and would not otherwise use.

d. Closing the project will cost the company an additional $3,000 in the final year of the project
Answer #1

Here,

Cost of Capital =10%

Initial cash outlay = $10000

This investment generate cash flow of $ 2000 per year for 10 year. And Present Value of Cash Flow =

= 2000* PVIFA(10,10%)

= 2000*6.1446

= $12289

Now, Net Present Value

= Present value of cash flow- Initial outlay

= 12289-10000

= $2289

Company have to invest in such project because its give positive NPV of $2289.

Answer a) Project risk

Risk is affect the future cash flow in calculation of net present value. If risk is higher than it can reduce the cash flow of project.

Answer b) Maintenance cost of $500

Maintenance Cost is outflow of cash its reduce the NPV of project

Here, PV of cost of maintenance = $500*PVIFA (10,10%)

= 500*6.1446

= $3072

NPV = 12289-10000-3072 = - 783

Answer c) Cost of $5000 of Equipment.

It was included initial outflow of cash, which reduce the NPV of project

Answer d) cost of closing the project

Cost of closing the project increase the value of cash outlay of project.

PV of Cost of closing the project = 3000*PVIF(10,10%)

= 3000*0.3855

= $1156.50

Now NPV = 12289-10000-1156.501= $1132.50